Casscatanzar84:

The expression 113.47 - (43.72 - 26.9) represents A. The expression 113.47 - (26.9 - 43.72) represents B.

darkknight:

What is the problem?

darkknight:

Question associated with this?

TheSmartOne:

The question is The numerical value of A - B is . The numerical value of B - A is .

darkknight:

@casscatanzar84, What do you think/

jhonyy9:

there are 3 numbers 113,47 43,72 26,9 the first expression is A and the second is B so what may be the question here ? for example what is true ? A > B A = B A < B

jhonyy9:

any idea ?

Shimansk123:

Find out The numerical value of A - B  and the numerical value of B - A. To prove As given The expression 113.47 - (43.72 - 26.9) represents A. The expression 113.47 - (26.9 - 43.72) represents B. Thus A - B =  113.47 - (43.72 - 26.9)  - ( 113.47 - (26.9 - 43.72)) First solving the bracket terms. A - B =  113.47 - (43.72 - 26.9)  -  113.47  + (26.9 - 43.72)         = 113.47 - 16.82 - 113.47 - 16.82         =  113.47  - 113.47 - 16.82 - 16.82         = -33.64 Therefore the value of A- B is -33.64. Thus B - A = 113.47 - (26.9 - 43.72) - (113.47 - (43.72 - 26.9)) First solving the bracket terms. B - A = 113.47 - (26.9 - 43.72) - 113.47 + (43.72 - 26.9)         =  113.47 + 16.82 - 113.47 + 16.82         = 33.64 Therefore the value of the A - B is -33.64 and B - A is 33.64